OpenStudy (anonymous):

Show that cot(θ) = -tan(θ+π/2).

OpenStudy (anonymous):

are you familiar with the formula for combined tan of two angles? \[Tan (A+B)= \frac{Tan A +Tan B}{1-TanA*TAnB}\]

OpenStudy (anonymous):

I am but I'm rather confused with the negative in front of the tan.

OpenStudy (anonymous):

it will be obtained in due time

OpenStudy (anonymous):

tans(A+90) = Tan A +Tan90 / 1-TanA*Tan90

OpenStudy (anonymous):

what is the value of Tan 90?

OpenStudy (anonymous):

I know that equals 1.

OpenStudy (anonymous):

I got at far as you gotten now but I'm just stuck right there on how I could get cot(θ).

OpenStudy (anonymous):

\[Tan(\theta+90) = \frac{Tan \theta+ Tan90}{1-Tan \theta*Tan 90}\] Taking out Tan90 as a common term \[\frac{\frac{Tan \theta}{Tan90}+1}{\frac{1}{Tan90}- Tan \theta}\]

OpenStudy (anonymous):

Now Tan 90 = infinity and 1/ infinity = 0 \[\frac{0+1}{0-Tan \theta} = \frac{-1}{Tan \theta} = - Cot \Theta\]

OpenStudy (anonymous):

\[Tan (\theta +90)= - Cot \theta \] Thus \[Cot \theta = - Tan (\theta +90) \]

OpenStudy (anonymous):

any doubts?

OpenStudy (anonymous):

Wow no doubts at all. Haha, thank you so much for the clarifying everything for me. :)

OpenStudy (anonymous):

Its been a pleasure :) take care

OpenStudy (anonymous):

I doubt most math teachers will take that. Division by zero isn't allowed, but neither is division by infinity.

OpenStudy (anonymous):

You might approach it like this.\[ -\tan(θ+π/2)=\frac{-\sin(θ+π/2)}{\cos(θ+π/2)}=\frac{-(-\cos \theta)}{\sin \theta}=\cot \theta\]

OpenStudy (anonymous):

Oh o: That works too and seems a lot shorter.

OpenStudy (anonymous):

No sweat. Do math every day.

OpenStudy (anonymous):

Thanks. :)